40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 10 15:54:14 2015, Steve, Update, VAC, Vacuum comp. rebooted afterReb061015.png
    Reply  Thu Jun 11 19:40:59 2015, Koji, Update, VAC, Vacuum comp. rebooted 
       Reply  Fri Jun 12 08:40:17 2015, Steve, Update, VAC, Vacuum comp. rebooted vacMonNoGauges.pngprepReboot.png
Message ID: 11353     Entry time: Thu Jun 11 19:40:59 2015     In reply to: 11352     Reply to this: 11354
Author: Koji 
Type: Update 
Category: VAC 
Subject: Vacuum comp. rebooted 

The serial connections to the vacuum gauges were recovered by rebooting c1vac1 and c1vac2.

Steve claimed that the vacuum screen had showed "NO COMM" at the vacuum pressure values.
The epics connection to c1vac was fine. We could logged in to c1vac1 with telnet too although c1vac2 had no response.

After some inspection, we decided to reboot the slow machines. Steve manually XXXed YYY valves (to be described)
to prepare for any possible unwanted switching. Initially Koji thought only c1vac2 can be rebooted. But it was wrong.
If the reset button is pushed, all of the modules on the same crate is reset. So everything was reset. After ~3min we still
don't have the connection to c1vac1 restored. We decided to another reboot. This time I pushed c1vac1 reset button.
After waiting about two minutes, the ADCs started to show green lights and the switch box started scanning.
We recovered the telnet connection to c1vac1 and epics functions. c1vac2 is still note responding to telnet, and
the values associated with c1vac2 are still blank.

Steve restored the valves and everything was back to normal.

ELOG V3.1.3-